OpenStudy (anonymous):

find the value of m such that the function y=e^mx is asolution of y`+2y=0

OpenStudy (anonymous):

Just plug in the function, like this: (e^mx)' = me^mx me^mx + 2e^mx = 0 since e^mx is never 0, we got: m + 2 = 0 therefore m = -1.

OpenStudy (kbee):

alieeeeeen true though supposed to be m = -2